the perimeter of a triangle is 73cm. the triangle is isosceles now, but if you lengthened the base by 5 cm and shortened each side by 6cm it would be equilateral. find the base of the original triangle.
Answer by ankor@dixie-net.com(22740)

the perimeter of a triangle is 73cm. the triangle is isosceles now, but if you lengthened the base by 5 cm and shortened each side by 6cm it would be equilateral. find the base of the original triangle.
:
let b = the base of the original triangle
let s = two equal sides
:
"the perimeter of a triangle is 73cm."
b + 2s = 73
" the triangle is isosceles now, but if you lengthened the base by 5 cm and shortened each side by 6cm it would be equilateral."
(sides would be equal)
(b+5) = (s-6)
b - s = -6 - 5
b - s = -11
:
Use elimination with these two equations
b + 2s = 73
b - s = -11
-------------Subtraction eliminates b, find s
0 + 3s = 84
s = 84/3
s = 28 cm, the equal side of the original triangle
:
" find the base of the original triangle."
Use b + 2s = 73
b + 2(28) = 73
b = 73 - 56
b = 17 cm the original base
:
:
:
Confirm this
"lengthened the base by 5 cm and shortened each side by 6cm it would be equilateral. "
17 + 5 = 28 - 6
22 = 22